# Large-scale distributed synchronization systems, using a cancel-on-completion redundancy mechanism

### v1

@@ 7 Main results
Proof. (i) At least one proper ML-FP exists by Theorem 7.2. Let $v_{min}\geq 0$ be the infimum of speeds, for which a proper ML-FP exists. By Assumption  3.1(ii), upon an arrival of a job, uniformly on the mutual positions of the particles it selects, the expected total (positive) displacement of the selected particles is bounded away from 00 by a constant $\delta>0$. Then, a proper ML-FP speed must be at least $v\geq\lambda\delta$. Therefore, $v_{min}>0$. Also, upon an arrival of a job, uniformly on the mutual positions of the particles it selects, the expected total (positive) displacement of the selected particles is upper bounded by the expected total size of all components, which is uniformly bounded by a finite $\delta^{\prime}$. Therefore, $v_{max}\leq\lambda\delta^{\prime}<\infty$. There must be ML-FPs for both speeds $v_{min}$ and $v_{max}$, by Lemma 5.12 and Theorem 7.1.
@@
Now, if ML-FPs exist for speeds $v$ and $\hat{v}$, with $v<\hat{v}$, then the ML-FP exists for any speed $v^{\prime}\in(v,\hat{v})$. Indeed, if we consider the point where the former two intersect, then there exists a unique DE-FP solution with speed $v^{\prime}$ with proper left tail, which contains that intersection point; both to the left and to the right of the intersection point, this solution is sandwiched between the former two ML-FPs, and therefore has to be proper on the right as well.
@@
The proof of Lemma 7.5 follows by coupling and monotonicity, if WLOG we assume that the initial location of the right boundary is at 00.
@@
Proof. Let us fix a small $\delta>0$, and assume WLOG that the initial state $x(0)$ of ML $x(\cdot)$ is such that $x_{0}(0)=1-\delta$. Then $x(\cdot)$ dominates ML $x^{\ell}(\cdot)$ with initial state $x^{\ell}(0)=(1-\delta)x^{\emptyset}\leq x(0)$; this initial state is such that mass $\delta$ is located at $\infty$- ∞ and mass $1-\delta$ is located at 00. We see that the velocity of $x^{\ell}(\cdot)$, if we only look at $\nu$-quantiles with $\nu\in(\delta,1)$, is same as the velocity of the ML starting from $x^{\emptyset}$ in the $(0,\delta)$-reduced system. Now note that in the $(0,\delta)$-reduced system the job class arrival rates $\{\hat{\sigma}_{j}\}$ are only slightly different from $\{\sigma_{j}\}$, and therefore $\{\tilde{\sigma}_{j}\}\leq\{\hat{\sigma}_{j}\}$. Applying Lemma 7.5, we see that the velocity of $x^{\ell}(\cdot)$, limited to $\nu$-quantiles with $\nu\in(\delta,1)$, is at least some $\tilde{v}<v$, which can be made arbitrarily close to $v$ by making $\delta$ small. Recalling that $x(\cdot)$ dominates $x^{\ell}(\cdot)$, we conclude that the velocity of $x(\cdot)$ is lower bounded by $v$. $\Box$
@@
Proof. Let us fix a small $\epsilon>0$. Consider the ML-FP $x^{*}$ such that $x^{*}_{0}=\epsilon$. Then, $x^{*}$ in the interval $w\leq 0$ is the unique ML-FP for the $(\epsilon,0)$-reduced system with the right boundary at 00. Since $\epsilon$ is small, the job class arrival rates $\{\hat{\sigma}_{j}\}$ in the reduced system are close to the job class arrival rates $\{\sigma_{j}\}$ in the original system. Since all $\sigma_{j}$ are positive (by Assumption 7.7), in the reduced system we can slightly reduce the speed from $v$ to $\tilde{v}$, and reduce $\lambda$ by the same factor to $\tilde{\lambda}=\lambda(\tilde{v}/v)$, so that the job class arrival rates $\{\tilde{\sigma}_{j}=\hat{\sigma}_{j}(\tilde{v}/v)\}$ in the reduced system are such that $\{\tilde{\sigma}_{j}\}<\{\sigma_{j}\}$. Note that $\tilde{v}$ can be arbitrarily close to $v$ if $\epsilon$ is small enough. Application of Theorem 7.6 completes the proof. $\Box$
@@
(i) If free ML-FP speed range is $[v_{min},v_{max}]$, then for the anchored left-regulated system with frame $[0,\infty)$, for any $v>v_{max}$, there exists a unique proper ML-FP (which is then equal to the ML-MFP), and Condition 6.5 holds.
@@
(iii) The existence of the limiting ML-FP $x^{*,max}$ with specified properties easily follows from Lemma 5.12. Consider $x^{*,max}$ with the left boundary at 00, and $x^{*,max,free}$ intersecting $x^{*,max}$ at $w=0$ (i.e., $x^{*,max,free}_{0}=x^{*,max}_{0}$). The specified family $\tilde{x}^{*}$ of left-regulated ML-FP is obtained by considering the family of DE-FP with the left boundary moving left to $\infty$- ∞, and intersecting $x^{*,max,free}$ and $x^{*,max}$ at the same point at $w=0$. The difference of inverses $\hat{\phi}_{u}-\tilde{\phi}_{u}$ is non-increasing for any two ML-FP in this family, including $x^{*,max,free}$ and $x^{*,max}$ as its extreme points, is by Theorem 7.4(ii). Then any $\tilde{x}^{*}$ is sandwiched between $x^{*,max,free}$ and $x^{*,max}$ both in $(-\infty,0]$ and in $[0,\infty)$, and therefore each $\tilde{x}^{*}$ is proper.
@@
Suppose Assumptions 3.2, 3.3 and 3.4 hold. If free ML-FP speed range is $[v_{min},v_{max}]$, then SSAI holds for the left-regulated systems with frame $[0,\infty)$ for speeds $v>v_{max}$, which correspond to ML-FP loads within $[0,\bar{\rho})$, $\bar{\rho}\leq 1$. Specifically, for each $v>v_{max}$ the $n$-th system is stable (for all sufficiently large $n$) and the SSAI holds, with the limiting distribution concentrated on the ML-MFP $x^{*,v}$ (which is the unique left-regulated proper ML-FP for this speed), and load $\rho(v)=x^{*,v}_{0}\uparrow\bar{\rho}$ as $v\downarrow v_{\max}$. (Note that if, in addition, Condition 6.5 holds for all $v$, then, by Theorem 7.11(v), $\bar{\rho}=1$, and therefore SSAI-FRL holds.)
Proof of Theorem 7.12. By Theorem 7.11, there is the family of unique, left-regulated ML-FP (=ML-MFP) $x^{*,v}$, with load $\rho(v)=x^{*,v}_{0}$ continuously strictly increasing from 00 to $\bar{\rho}$ as $v$ decreases from +$\infty$+ ∞ to $v_{max}$; and Condition 6.5 holds for any $v>v_{max}$. Denote by $v(\rho)$ the inverse of $\rho(v)$.
@@
(i) If $v_{min}=v_{max}=v_{*}$, then $\bar{\rho}$ in Theorem 7.12 is the supremum of the loads $x^{*,v}_{0}$ of all proper left-regulated ML-MFP $x^{*,v}$ in the frame $[0,\infty)$, under different speeds $v$.
@@
Suppose Assumptions 3.2, 3.3 and 3.4 hold. If free ML-FP speed range is $[v_{min},v_{max}]$, then SSAI holds for the right-regulated systems with frame $(-\infty,0]$ for speeds $v<v_{min}$. Specifically, for each $v<v_{min}$ the $n$-th system is stable (for all sufficiently large $n$) and the SSAI holds, with the limiting distribution concentrated on the ML-MFP $x^{*,v}$ (which is the unique right-regulated proper ML-FP for this speed). Furthermore, as $v\uparrow v_{min}$, (a shifted version of) $x^{*,v}$ converges to the unique proper free ML-FP for speed $v_{min}$.
@@
(i) Suppose the centered free process $\mathring{x}^{n}(\cdot)$ is positive recurrent, and therefore the average velocity $v_{n}$ of unanchored particle system is well-defined – see (19)-(20). (Positive recurrence holds, for example, if, in addition, Assumption 6.2 holds – see Theorem 6.4.) Then, $\liminf_{n}v_{n}\geq v_{min}$ and $\limsup_{n}v_{n}\leq v_{max}$, where $[v_{min},v_{max}]$ is the free proper ML-FP speed range.
@@
[…]
@@
Consider now the $n$-th pre-limit unanchored process $x^{n}(\cdot)$ (with $n$ particles) in the stationary regime, in the sense that the centered process $\mathring{x}^{n}(\cdot)$ is in the stationary regime. WLOG, let the $\delta$-quantile of the process (random) initial state $x^{n}(0)$ be located exactly at 00. Consider $\nu$-quantile of $x^{n}(0)$, with $\nu\in(\delta,1)$. The $\nu$-quantile’s expected distance from the $\delta$-quantile is uniformly bounded in $n$. (Here we use the fact that, by Theorem 6.4, $\mathbb{E}\Phi_{1}(\mathring{x}^{n}(\infty))$ is finite, uniformly bounded.) If we consider a different initial state, with $\delta$-fraction of particles moved back to $\infty$- ∞ and the remaining $(1-\delta)$-fraction of particles moved back to $\delta$-quantile, then, by monotonicity, we obtain a lower bounding process which is dominated by $x^{n}(t)$ at all times $t$. But, as $n\to\infty$, the ML for this lower bounding process is exactly $x^{\ell}(\cdot)$, whose velocity (restricted to $\nu$-quantiles with $\nu\in(\delta,1)$) is at least $\tilde{v}$. Using this, (20) and (19), we can write
@@
To construct a lower bounding system, consider a class, satisfying Assumption 3.3. Then, WLOG, there is a class $j$ such that there are $m_{j}$ components, $0\leq m_{j}<k_{j}$, which are exactly 00, and the remaining $d_{j}-m_{j}$ components have random sizes which are between $\epsilon>0$ and $\epsilon_{2}>\epsilon$. We define a lower bounding system by replacing class $j$ by two classes as follows: the first one occurs with high probability $11-\delta_{1}<$ 1 and has the same component size distribution $F_{j}$ as class $j$; the second one occurs with small probability $\delta_{1}>0$, and the component size distribution is such that we, first, take a realization of sizes according to $F_{j}$ and then replace each non-zero component size realization $s$ by realization of the uniform distribution in $[0,s]$. The second of the new classes (of the lower bounding system) satisfies conditions of Lemma 6.3. Therefore, the lower bounding system satisfies Assumption 6.2. Now, as $\delta_{1}\downarrow 0$ along some sequence, conditions of Theorem 7.10 are satisfied, and so we can choose a lower bounding system with $v_{min}$ being arbitrarily close to that of the original system. $\Box$
@@ 8 Proofs
Since a stationary distribution is invariant w.r.t. simultaneous scaling of speed $v$ and $\lambda^{n}$, let us assume that $v=1$ WLOG. Then the evolution of $p^{n}(t)$ is as follows. Note that here the particle locations $w_{i}(t)$ may be negative as well as non-negative. Point $z(t)$ serves as a regulation boundary, which evolves in time; $z(t)\leq 0$ at all times. Denote by $I_{z}(t)$ the (possibly empty) subset of particles, which are located exactly at the regulation boundary $z(t)$ at time $t$. Between the times of the particle jumps (job arrivals), the boundary $z(t)$ moves right at the constant non-negative speed $(n-|I_{z}(t)|)/n$; the particles in $I_{z}(t)$ (those located exactly at the boundary $z(t)$) stay at the boundary and therefore move with it to the right at constant non-negative speed $(n-|I_{z}(t)|)/n$; each particle that is not at $z(t)$ (i.e. not within subset $I_{z}(t)$) moves to the left at the constant non-positive speed $-|I_{z}(t)|/n$; when a particle “hits” boundary $z(t)$, it joins the set $I_{z}(t)$. (The average of the particle locations stays at 00, as it should by the $p(t)$ definition.) It is easy to see that, for any $\ell\geq 1$, $\Phi_{\ell}(p^{n}(t))$ is absolutely continuous with the derivative (existing almost everywhere) $(d/dt)\Phi_{\ell}(p^{n}(t))\leq 0$, and moreover $(d/dt)\Phi_{\ell}(p(t))<0$ as long as $0<|I_{z}(t)|<n$. At a time when one or more particles jump (upon a job arrival), the following occurs. Let $\kappa_{i}(t)$ be the amount of new workload added to server $i$ by the job arriving at time $t$ (which may be non-zero only for the servers within the selection set of the job). Then, particle $i$ jump size at $t$ is $\zeta_{i}=\kappa_{i}(t)-\sum_{s}\kappa_{s}(t)/n$ (which may be positive or negative); the point $z(t)$ jumps (left) by $-\sum_{s}\kappa_{s}(t)/n$.
@@
$\bar{\zeta}^{(\mathring{x}^{n})}(w)-\bar{\zeta}^{(\mathring{x}^{n})}(0)$ is non-increasing with value 00 at 00, and $\mathring{x}^{n}\in\mathring{{\cal X}}_{1}$ (and then $\int_{-\infty}^{\infty}wd[-\mathring{x}^{n}_{w}]=0$).
@@
The definition of $\bar{\zeta}_{i}=\bar{\zeta}_{i}^{(\mathring{x}^{n})}$ in (30) can be interpreted as follows: $\bar{\zeta}_{i}$ is the expected actual jump size $\mathbb{E}[\kappa_{i}|i\in S]$ of particle $i$, conditioned on it being selected by a job, multiplied by constant $[n{\mathbb{P}}\{i\in S\}]$ (which is close to the constant $\sum_{j}\pi_{j}d_{j}$ for all large $n$), and then centered by the total expected particles’ displacement $\mathbb{E}\sum_{s}\kappa_{s}$ due to a job arrival.
@@
(If not, since both $a^{+}_{\nu,n}/c_{n}$ and $a^{-}_{\nu,n}/c_{n}$ are bounded away from 00, $\nu$ could not be small.) Let us fix any $\delta_{1}\in(0,1/2)$, and then choose a corresponding $\nu$.
@@
Indeed, when a job arrival selects both $i^{-}$ and $i^{+}$, then, conditioned on this event, by monotonicity, $i^{-}$ receives at least as much average new workload as $i^{+}$. Also, clearly, the probability that both $i^{-}$ and $i^{+}$ are selected by an arriving job, under the condition that $i^{-}$ (respectively, $i^{+}$) is selected vanishes as $n\to\infty$. Thus, to show (38), is suffices to consider two probability spaces, $E^{-}$ and $E^{+}$, obtained by conditioning on the two non-intersecting events, when, respectively, $i^{-}$ (but not $i^{+}$) and $i^{-}$ (but not $i^{+}$) is selected by a job. We can couple the constructions of the spaces $E^{-}$ and $E^{+}$, so that, w.p.1, $i^{-}$ (on $E^{-}$) and $i^{+}$ (on $E^{+}$) receive equal component sizes, while all other selected servers and their component sizes are equal. At least one of the following two conditions holds for $\mathring{x}^{n}$: (a) at least $[(1-2\delta_{1})n/2]/2$ particles are located in $[0,a^{+}_{\nu,n})$ or (b) at least $[(1-2\delta_{1})n/2]/2$ particles are located in $(a^{-}_{\nu,n},0]$. In the case (a), consider an arrival of a job of class $j$, satisfying Assumption 3.4. Then, with probability bounded away from 00 all other selected particles are in $[0,a^{+}_{\nu,n})$; and then, Assumption 3.4 easily implies (38), because, with bounded away from 00 probability, $i^{-}$ will receive new workload equal to its component size, while $i^{+}$ will receive a strictly smaller (by a bounded away from 00 quantity) new workload. In the case (b), consider an arrival of a job of class $j$, satisfying Assumption 3.3. With probability bounded away from 00 all other selected particles are in $(a^{-}_{\nu,n},0]$; and then, Assumption 3.3 easily implies (38), because, with bounded away from 00 probability, $i^{-}$ will make some positive (bounded away from 00) jump forward, while $i^{+}$ will not jump. This completes the proof of (38).
@@
Consider the class $j$ as in the lemma statement. WLOG, we can assume that this class is such that $m_{j}$ components are zero and the remaining $d_{j}-m_{j}$ components distribution is such that those components are i.i.d., uniformly distributed in $[0,\epsilon]^{d_{j}-m_{j}}$. (We can always split the original class $j$ into two, with one of the new classes satisfying the condition above.) Fix any $C>0$ as in the definition of a small set. Fix arbitrarily small $\delta_{2}>0$. Then, uniformly in all initial states with norm at most $C$, with strictly positive probability, the system gets into a state with norm (= max distance between particles) at most $\delta_{2}$. Now consider the special system state such that all particles are exactly co-located. Further, consider a random state, obtained by a sufficiently large, fixed number (say $n$) of arrivals of this class, such that each particle has non-zero component at least once, with the included additional condition that all generated non-zero components happen to fall into $[\epsilon/3,2\epsilon/3]$; the distribution of this random state – let us denote it $\mathcal{D}$ – is non-proper - the total measure is less than $1$. We claim the following: if $\delta_{2}$ is small enough, and the initial state norm is at most $\delta_{2}$, and we generate a new state exactly the same way as we generated $\mathcal{D}$, except the non-zero components are distributed in $[0,\epsilon]$ (i.e., without including the additional condition), then the distribution of new random state will dominate $\mathcal{D}$. The claim is true, because for each outcome in the special system, there exists an “equally probable” outcome in the actual system, obtained by slightly shifting, for each particle, its first the non-zero component size to compensate for the deviation of the initial state from special initial state. Therefore, the system satisfies Assumption 6.2. $\Box$
@@
[…]
the average drift of the particle at point $w$. It is continuous strictly decreasing at all $w$, except maybe a left boundary point (if any), where it may have a discontinuity – jump down. The average of $\bar{\psi}_{w}$ w.r.t. distribution $x$ is 00. This means that $\bar{\psi}_{w}<0$ for some $w$ sufficiently far on the right.
@@
If there is no left boundary, $\bar{\psi}_{w}$ is continuous, and there is a point $u$ such that $\bar{\psi}_{u}=0$. Let us choose this point for centering and consider Lyapunov function $(w-u)^{2}\wedge C_{2}$. Using analogous generator upper bound, we can show that $\mathbb{E}|X(\infty)-u|=\infty$ is impossible, because it would imply negative average drift of $[X(t)-u]^{2}\wedge C_{2}$ when $C_{2}$ is large. Therefore, $\mathbb{E}|X(\infty)-u|<\infty$, and $\bar{x}$ is well-defined finite. (A small subtlety here. If $X(t)$ jumps so that $X(t+)>u+\sqrt{C_{2}}$, then replacing $[X(t+)-u]^{2}\wedge C_{2}=C_{2}$ by $[X(t+)-u]^{2}$ in the drift estimate, makes the estimate larger.)

### v2

@@ 7 Main results
Proof. (i) At least one proper ML-FP exists by Theorem 7.2. Let $v_{min}\geq 0$ be the infimum of speeds, for which a proper ML-FP exists. By Assumption  3.1(ii), upon an arrival of a job, uniformly on the mutual positions of the particles it selects, the expected total (positive) displacement of the selected particles is bounded away from 0 by a constant $\delta>0$. Then, a proper ML-FP speed must be at least $v\geq\lambda\delta$. Therefore, $v_{min}>0$. Also, upon an arrival of a job, uniformly on the mutual positions of the particles it selects, the expected total (positive) displacement of the selected particles is upper bounded by the expected total size of all components, which is uniformly bounded by a finite $\delta^{\prime}$. Therefore, $v_{max}\leq\lambda\delta^{\prime}<\infty$. There must be ML-FPs for both speeds $v_{min}$ and $v_{max}$, by Lemma 5.12 and Theorem 7.1.
@@
[…]
@@
The proof of Lemma 7.5 follows by coupling and monotonicity, if WLOG we assume that the initial location of the right boundary is at 0.
@@
Proof. Let us fix a small $\delta>0$, and assume WLOG that the initial state $x(0)$ of ML $x(\cdot)$ is such that $x_{0}(0)=1-\delta$. Then $x(\cdot)$ dominates ML $x^{\ell}(\cdot)$ with initial state $x^{\ell}(0)=(1-\delta)x^{\emptyset}\leq x(0)$; this initial state is such that mass $\delta$ is located at -$\infty$ and mass $1-\delta$ is located at 0. We see that the velocity of $x^{\ell}(\cdot)$, if we only look at $\nu$-quantiles with $\nu\in(\delta,1)$, is same as the velocity of the ML starting from $x^{\emptyset}$ in the $(0,\delta)$-reduced system. Now note that in the $(0,\delta)$-reduced system the job class arrival rates $\{\hat{\sigma}_{j}\}$ are only slightly different from $\{\sigma_{j}\}$, and therefore $\{\tilde{\sigma}_{j}\}\leq\{\hat{\sigma}_{j}\}$. Applying Lemma 7.5, we see that the velocity of $x^{\ell}(\cdot)$, limited to $\nu$-quantiles with $\nu\in(\delta,1)$, is at least some $\tilde{v}<v$, which can be made arbitrarily close to $v$ by making $\delta$ small. Recalling that $x(\cdot)$ dominates $x^{\ell}(\cdot)$, we conclude that the velocity of $x(\cdot)$ is lower bounded by $v$. $\Box$
@@
Proof. Let us fix a small $\epsilon>0$. Consider the ML-FP $x^{*}$ such that $x^{*}_{0}=\epsilon$. Then, $x^{*}$ in the interval $w\leq 0$ is the unique ML-FP for the $(\epsilon,0)$-reduced system with the right boundary at 0. Since $\epsilon$ is small, the job class arrival rates $\{\hat{\sigma}_{j}\}$ in the reduced system are close to the job class arrival rates $\{\sigma_{j}\}$ in the original system. Since all $\sigma_{j}$ are positive (by Assumption 7.7), in the reduced system we can slightly reduce the speed from $v$ to $\tilde{v}$, and reduce $\lambda$ by the same factor to $\tilde{\lambda}=\lambda(\tilde{v}/v)$, so that the job class arrival rates $\{\tilde{\sigma}_{j}=\hat{\sigma}_{j}(\tilde{v}/v)\}$ in the reduced system are such that $\{\tilde{\sigma}_{j}\}<\{\sigma_{j}\}$. Note that $\tilde{v}$ can be arbitrarily close to $v$ if $\epsilon$ is small enough. Application of Theorem 7.6 completes the proof. $\Box$
@@
(i) If free ML-FP speed range is $v_{min},v_{max}]$, then for the anchored left-regulated system with frame $0,\infty)$, for any $v>v_{max}$, there exists a unique proper ML-FP (which is then equal to the ML-MFP), and Condition 6.5 holds.
@@
(iii) The existence of the limiting ML-FP $x^{*,max}$ with specified properties easily follows from Lemma 5.12. Consider $x^{*,max}$ with the left boundary at 0, and $x^{*,max,free}$ intersecting $x^{*,max}$ at $w=0$ (i.e., $x^{*,max,free}_{0}=x^{*,max}_{0}$). The specified family $\tilde{x}^{*}$ of left-regulated ML-FP is obtained by considering the family of DE-FP with the left boundary moving left to -$\infty$, and intersecting $x^{*,max,free}$ and $x^{*,max}$ at the same point at $w=0$. The difference of inverses $\hat{\phi}_{u}-\tilde{\phi}_{u}$ is non-increasing for any two ML-FP in this family, including $x^{*,max,free}$ and $x^{*,max}$ as its extreme points, is by Theorem 7.4(ii). Then any $\tilde{x}^{*}$ is sandwiched between $x^{*,max,free}$ and $x^{*,max}$ both in $(-\infty,0]$ and in $0,\infty)$, and therefore each $\tilde{x}^{*}$ is proper.
@@
Suppose Assumptions 3.2, 3.3 and 3.4 hold. If free ML-FP speed range is $v_{min},v_{max}]$, then SSAI holds for the left-regulated systems with frame $0,\infty)$ for speeds $v>v_{max}$, which correspond to ML-FP loads within $0,\bar{\rho})$, $\bar{\rho}\leq 1$. Specifically, for each $v>v_{max}$ the $n$-th system is stable (for all sufficiently large $n$) and the SSAI holds, with the limiting distribution concentrated on the ML-MFP $x^{*,v}$ (which is the unique left-regulated proper ML-FP for this speed), and load $\rho(v)=x^{*,v}_{0}\uparrow\bar{\rho}$ as $v\downarrow v_{\max}$. (Note that if, in addition, Condition 6.5 holds for all $v$, then, by Theorem 7.11(v), $\bar{\rho}=1$, and therefore SSAI-FRL holds.)
Proof of Theorem 7.12. By Theorem 7.11, there is the family of unique, left-regulated ML-FP (=ML-MFP) $x^{*,v}$, with load $\rho(v)=x^{*,v}_{0}$ continuously strictly increasing from 0 to $\bar{\rho}$ as $v$ decreases from $\infty$ to $v_{max}$; and Condition 6.5 holds for any $v>v_{max}$. Denote by $v(\rho)$ the inverse of $\rho(v)$.
@@
(i) If $v_{min}=v_{max}=v_{*}$, then $\bar{\rho}$ in Theorem 7.12 is the supremum of the loads $x^{*,v}_{0}$ of all proper left-regulated ML-MFP $x^{*,v}$ in the frame $0,\infty)$, under different speeds $v$.
@@
Suppose Assumptions 3.2, 3.3 and 3.4 hold. If free ML-FP speed range is $v_{min},v_{max}]$, then SSAI holds for the right-regulated systems with frame $(-\infty,0]$ for speeds $v<v_{min}$. Specifically, for each $v<v_{min}$ the $n$-th system is stable (for all sufficiently large $n$) and the SSAI holds, with the limiting distribution concentrated on the ML-MFP $x^{*,v}$ (which is the unique right-regulated proper ML-FP for this speed). Furthermore, as $v\uparrow v_{min}$, (a shifted version of) $x^{*,v}$ converges to the unique proper free ML-FP for speed $v_{min}$.
@@
(i) Suppose the centered free process $\mathring{x}^{n}(\cdot)$ is positive recurrent, and therefore the average velocity $v_{n}$ of unanchored particle system is well-defined – see (19)-(20). (Positive recurrence holds, for example, if, in addition, Assumption 6.2 holds – see Theorem 6.4.) Then, $\liminf_{n}v_{n}\geq v_{min}$ and $\limsup_{n}v_{n}\leq v_{max}$, where $v_{min},v_{max}]$ is the free proper ML-FP speed range.
@@
$v_{min,n}=\sup\{v~|~\mbox{$\forall\nu\in(0,1),~\liminf_{t\to\infty}\Delta^{n}(t;\nu)/t\geq v$, w.p.1}\}$,
$v_{max,n}=\inf\{v~|~\mbox{$\forall\nu\in(0,1),~\limsup_{t\to\infty}\Delta^{n}(t;\nu)/t\leq v$, w.p.1}\}$.
@@
Consider now the $n$-th pre-limit unanchored process $x^{n}(\cdot)$ (with $n$ particles) in the stationary regime, in the sense that the centered process $\mathring{x}^{n}(\cdot)$ is in the stationary regime. WLOG, let the $\delta$-quantile of the process (random) initial state $x^{n}(0)$ be located exactly at 0. Consider $\nu$-quantile of $x^{n}(0)$, with $\nu\in(\delta,1)$. The $\nu$-quantile’s expected distance from the $\delta$-quantile is uniformly bounded in $n$. (Here we use the fact that, by Theorem 6.4, $\mathbb{E}\Phi_{1}(\mathring{x}^{n}(\infty))$ is finite, uniformly bounded.) If we consider a different initial state, with $\delta$-fraction of particles moved back to -$\infty$ and the remaining $(1-\delta)$-fraction of particles moved back to $\delta$-quantile, then, by monotonicity, we obtain a lower bounding process which is dominated by $x^{n}(t)$ at all times $t$. But, as $n\to\infty$, the ML for this lower bounding process is exactly $x^{\ell}(\cdot)$, whose velocity (restricted to $\nu$-quantiles with $\nu\in(\delta,1)$) is at least $\tilde{v}$. Using this, (20) and (19), we can write
@@
To construct a lower bounding system, consider a class, satisfying Assumption 3.3. Then, WLOG, there is a class $j$ such that there are $m_{j}$ components, $0\leq m_{j}<k_{j}$, which are exactly 0, and the remaining $d_{j}-m_{j}$ components have random sizes which are between $\epsilon>0$ and $\epsilon_{2}>\epsilon$. We define a lower bounding system by replacing class $j$ by two classes as follows: the first one occurs with high probability $1-\delta_{1}<1$ and has the same component size distribution $F_{j}$ as class $j$; the second one occurs with small probability $\delta_{1}>0$, and the component size distribution is such that we, first, take a realization of sizes according to $F_{j}$ and then replace each non-zero component size realization $s$ by realization of the uniform distribution in $[0,s]$. The second of the new classes (of the lower bounding system) satisfies conditions of Lemma 6.3. Therefore, the lower bounding system satisfies Assumption 6.2. Now, as $\delta_{1}\downarrow 0$ along some sequence, conditions of Theorem 7.10 are satisfied, and so we can choose a lower bounding system with $v_{min}$ being arbitrarily close to that of the original system. $\Box$
@@ 8 Proofs
Since a stationary distribution is invariant w.r.t. simultaneous scaling of speed $v$ and $\lambda^{n}$, let us assume that $v=1$ WLOG. Then the evolution of $p^{n}(t)$ is as follows. Note that here the particle locations $w_{i}(t)$ may be negative as well as non-negative. Point $z(t)$ serves as a regulation boundary, which evolves in time; $z(t)\leq 0$ at all times. Denote by $I_{z}(t)$ the (possibly empty) subset of particles, which are located exactly at the regulation boundary $z(t)$ at time $t$. Between the times of the particle jumps (job arrivals), the boundary $z(t)$ moves right at the constant non-negative speed $(n-|I_{z}(t)|)/n$; the particles in $I_{z}(t)$ (those located exactly at the boundary $z(t)$) stay at the boundary and therefore move with it to the right at constant non-negative speed $(n-|I_{z}(t)|)/n$; each particle that is not at $z(t)$ (i.e. not within subset $I_{z}(t)$) moves to the left at the constant non-positive speed $|I_{z}(t)|/n$; when a particle “hits” boundary $z(t)$, it joins the set $I_{z}(t)$. (The average of the particle locations stays at 0, as it should by the $p(t)$ definition.) It is easy to see that, for any $\ell\geq 1$, $\Phi_{\ell}(p^{n}(t))$ is absolutely continuous with the derivative (existing almost everywhere) $(d/dt)\Phi_{\ell}(p^{n}(t))\leq 0$, and moreover $(d/dt)\Phi_{\ell}(p(t))<0$ as long as $0<|I_{z}(t)|<n$. At a time when one or more particles jump (upon a job arrival), the following occurs. Let $\kappa_{i}(t)$ be the amount of new workload added to server $i$ by the job arriving at time $t$ (which may be non-zero only for the servers within the selection set of the job). Then, particle $i$ jump size at $t$ is $\zeta_{i}=\kappa_{i}(t)-\sum_{s}\kappa_{s}(t)/n$ (which may be positive or negative); the point $z(t)$ jumps (left) by $\sum_{s}\kappa_{s}(t)/n$.
@@
$\bar{\zeta}^{(\mathring{x}^{n})}(w)-\bar{\zeta}^{(\mathring{x}^{n})}(0)$ is non-increasing with value 0 at 0, and $\mathring{x}^{n}\in\mathring{{\cal X}}_{1}$ (and then $\int_{-\infty}^{\infty}wd[-\mathring{x}^{n}_{w}]=0$).
@@
The definition of $\bar{\zeta}_{i}=\bar{\zeta}_{i}^{(\mathring{x}^{n})}$ in (30) can be interpreted as follows: $\bar{\zeta}_{i}$ is the expected actual jump size $\mathbb{E}[\kappa_{i}|i\in S]$ of particle $i$, conditioned on it being selected by a job, multiplied by constant $n{\mathbb{P}}\{i\in S\}]$ (which is close to the constant $\sum_{j}\pi_{j}d_{j}$ for all large $n$), and then centered by the total expected particles’ displacement $\mathbb{E}\sum_{s}\kappa_{s}$ due to a job arrival.
@@
(If not, since both $a^{+}_{\nu,n}/c_{n}$ and $a^{-}_{\nu,n}/c_{n}$ are bounded away from 0, $\nu$ could not be small.) Let us fix any $\delta_{1}\in(0,1/2)$, and then choose a corresponding $\nu$.
@@
[…]
@@
Consider the class $j$ as in the lemma statement. WLOG, we can assume that this class is such that $m_{j}$ components are zero and the remaining $d_{j}-m_{j}$ components distribution is such that those components are i.i.d., uniformly distributed in $0,\epsilon]^{d_{j}-m_{j}}$. (We can always split the original class $j$ into two, with one of the new classes satisfying the condition above.) Fix any $C>0$ as in the definition of a small set. Fix arbitrarily small $\delta_{2}>0$. Then, uniformly in all initial states with norm at most $C$, with strictly positive probability, the system gets into a state with norm (= max distance between particles) at most $\delta_{2}$. Now consider the special system state such that all particles are exactly co-located. Further, consider a random state, obtained by a sufficiently large, fixed number (say $n$) of arrivals of this class, such that each particle has non-zero component at least once, with the included additional condition that all generated non-zero components happen to fall into $\epsilon/3,2\epsilon/3]$; the distribution of this random state – let us denote it $\mathcal{D}$ – is non-proper - the total measure is less than $1$. We claim the following: if $\delta_{2}$ is small enough, and the initial state norm is at most $\delta_{2}$, and we generate a new state exactly the same way as we generated $\mathcal{D}$, except the non-zero components are distributed in $0,\epsilon]$ (i.e., without including the additional condition), then the distribution of new random state will dominate $\mathcal{D}$. The claim is true, because for each outcome in the special system, there exists an “equally probable” outcome in the actual system, obtained by slightly shifting, for each particle, its first the non-zero component size to compensate for the deviation of the initial state from special initial state. Therefore, the system satisfies Assumption 6.2. $\Box$
@@
$\bar{\psi}_{w}=-v+\alpha\int_{0}^{\infty}\bar{J}^{(x,w)}(u)du,$
the average drift of the particle at point $w$. It is continuous strictly decreasing at all $w$, except maybe a left boundary point (if any), where it may have a discontinuity – jump down. The average of $\bar{\psi}_{w}$ w.r.t. distribution $x$ is 0. This means that $\bar{\psi}_{w}<0$ for some $w$ sufficiently far on the right.
@@
If there is no left boundary, $\bar{\psi}_{w}$ is continuous, and there is a point $u$ such that $\bar{\psi}_{u}=0$. Let us choose this point for centering and consider Lyapunov function $(w-u)^{2}\wedge C_{2}$. Using analogous generator upper bound, we can show that $\mathbb{E}|X(\infty)-u|=\infty$ is impossible, because it would imply negative average drift of $X(t)-u]^{2}\wedge C_{2}$ when $C_{2}$ is large. Therefore, $\mathbb{E}|X(\infty)-u|<\infty$, and $\bar{x}$ is well-defined finite. (A small subtlety here. If $X(t)$ jumps so that $X(t+)>u+\sqrt{C_{2}}$, then replacing $X(t+)-u]^{2}\wedge C_{2}=C_{2}$ by $X(t+)-u]^{2}$ in the drift estimate, makes the estimate larger.)